When Dennis asked Mr Wilson his age Mr Wilson refused to tell finally however he admitted that in 10 years he would be twice as old as he was 27 years ago how old is Mr Wilson now
Answer by ikleyn(52781)

.
The words translated to Math lead to this equation.


    W + 10 = 2*(W-27)


Solve it step by step


    W + 10 = 2W - 54

    10 + 54 = 2W - W

      64    = W.


ANSWER.  Mr. Wilson is 64 years years old now.

Solved.